# From Counterfactuals to Trees:Competitive Analysis of Model Extraction Attacks

### v1

@@ Appendix A Proofs
    Each subspace contains subtrees $f_{n_{1}}$ and $f_{n_{2}}$ with $n_{1},n_{2}\leq n$ split levels, respectively. By the inductive hypothesis, the number of queries for each subtree is respectively $O\left(\prod_{\begin{subarray}{c}i=1\\ s^{(n_{1})}_{i}\neq 0\end{subarray}}^{m}s^{(n_{1})}_{i}\right)$ and $O\left(\prod_{\begin{subarray}{c}i=1\\ s^{(n_{2})}_{i}\neq 0\end{subarray}}^{m}s^{(n_{2})}_{i}\right)$.
@@
    $O\left(\prod_{\begin{subarray}{c}i=1\\ s^{(n+1)}_{i}\neq 0\end{subarray}}^{m}s^{(n+1)}_{i}\right).$
Therefore, the query count for $f_{n}$ is $O\left(\prod_{\begin{subarray}{c}i=1\\ s^{(n)}_{i}\neq 0\end{subarray}}^{m}s^{(n)}_{i}\right)$. ∎
@@
$\max_{s^{(n)}_{1},\ldots,s^{(n)}_{p}}\prod_{i=1}^{p}s^{(n)}_{i}\quad\text{s.t.}\quad\sum_{i=1}^{p}s^{(n)}_{i}=n\quad\text{and}\quad s^{(n)}_{i}\geq 1\ \forall i\in\{1,\ldots,p\}.$
@@
$\max_{s^{(n)}_{1},\ldots,s^{(n)}_{p}}\sum_{i=1}^{p}\log(s^{(n)}_{i}).$
Applying the Karush-Kuhn-Tucker (KKT) conditions, we find that the maximum occurs when $s^{(n)}_{i}=\frac{n}{p}$ for all $i$. Substituting back, the worst-case complexity becomes $O\left(\left({n}/{p}\right)^{p}\right)$. ∎
@@
$\forall 1\leq j\leq m,\quad\sum_{p=1}^{j-1}s_{p}+1\leq i\leq\sum_{p=1}^{j}s_{p},\quad\alpha_{i}<\alpha_{i+1}.$
@@
    In the best-case scenario, where the decision tree follows a single branch (i.e., all splits occur along one feature), any algorithm, including the optimal one, would require at least $12n+$ 1 queries to reconstruct the tree (i.e. $Q_{opt}^{f}\geq 2n+1$). This includes one query for each split and additional queries to verify the leaf nodes.
    Conversely, in the worst-case scenario, such as a chessboard-like decision tree where splits are evenly distributed across multiple features, the TRA algorithm must explore all possible regions created by these splits. For a two-dimensional tree, this results in $s_{1}s_{2}+s_{1}+s_{2}+1$ queries, where $s_{1}$ and $s_{2}$ are the number of splits along each feature. Extending this to $m$ dimensions, the number of queries grows multiplicatively with the number of splits per feature, leading to:
@@
    $C_{TRA}^{(n,m)}=\text{sup}_{f\in\mathcal{F}}\left(\frac{Q_{TRA}^{f}}{Q_{opt}^{f}}\right)\leq\frac{\sum_{i=1}^{m}s_{i}\prod_{j=1}^{i-1}(s_{j}+1)+\prod_{j=1}^{m}(s_{j}+1)}{2n+1}$
@@
    Proof of Lower Bound:
@@
    Consider the following adversarial example: for each $1\leq i\leq n$ and $1\leq j\leq m$, let the dimension that $\alpha_{i}$ splits on be $j$, and set
    $\alpha_{i}=\begin{cases}\frac{i}{(s_{1}+1)},&\text{if }j=1,\\ \frac{i-\sum_{p=1}^{j-1}s_{p}}{2(s_{j}+1)}+\frac{1}{2}+\epsilon,&\text{otherwise},\end{cases}$
    where $\epsilon>0$. This adversarial example ensures that within any hyper-rectangle defined by split level boundaries, for $j=1,...,m-1$ if there are splits in both the $j$-th and $j+1$-th dimensions, then there exists a split in the $j$-th dimension that is closer to the center of the hyper-rectangle than any split in the $j+1$-th dimension.
    Therefore, in this adversarial example, TRA will always detect the splits of the $j$-th dimension before those of the $j+1$-th dimension. Therefore, the adversary can design a decision tree with a single branch (as illustrated in the right tree of Figure 6) that begins by splitting on the split levels in reverse (decreasing) order of dimensions (see a 2D example in Figure 6). For this specific example, TRA will require
@@
    $C_{TRA}^{(n,m)}\geq\frac{\sum_{i=1}^{m}s_{i}\prod_{j=1}^{i-1}(s_{j}+1)+\prod_{j=1}^{m}(s_{j}+1)}{2n+1}$
Hence,
@@
Figure 6: Adversarial example for TRA. The classes are $c_{1}$,$c_{2}$ and $c_{3}$. For simplicity, we denote $s_{1}=s^{(n)}_{1}$ and $s_{2}=s^{(n)}_{2}$. Here, the instance is dynamically built so that the number of queries required by TRA is $s_{2}(s_{1}+1)+s_{1}+(s_{1}+1)(s_{2}+1)$, whereas the optimal offline algorithm only needs $12n+$ 1, as shown in the right tree figure.
@@
We define a pure D&C-based algorithm as one that divides the input space (problem) into subspaces (sub-problems) based on counterfactual explanations and recursively continues this process within each subspace until no counterfactuals are found. This class of algorithms encompasses all types of querying strategies, such as selecting the top-left corner, bottom-right corner, or a random point within the input space, among others.
@@
    $1+\mathcal{Q}_{1}+\mathcal{Q}_{2}=(2n+1)C_{TRA}^{(n,m)}$
    given that $s_{1}^{(1)}+s_{1}^{(2)}+1=s_{1}$. Hence, the best competitive ratio $C^{(n,m)}_{D\&C}$ achievable by any D&C-based algorithm satisfies:
@@ B.1 Experimental Setup Details
For surrogate models that do not utilize the target model’s hyperparameters, we employ the default parameters provided by scikit-learn. Specifically for MLPs, we configure a scikit-learn MLP with two hidden layers, each consisting of 20 neurons, while keeping all other parameters at their default values.
@@
The anytime fidelity was calculated each 20 queries during all attacks execution, except for PathFinding which is not an anytime attack.
@@ B.2 Configuration of Surrogate-Based Attacks
We report in Figure 8 (respectively, Figure 9) the anytime performance of the CF (respectively, DualCF) model extraction attack against decision tree models, for the three considered types of surrogates and the two counterfactual oracles, on all considered datasets. More precisely, as depicted in Section 4.1, we run these two attacks using three different assumptions on adversarial knowledge, namely the hypothesis class of the target model, its hyperparameters, and none of them. In the first case, the adversary trains a decision tree surrogate with default parameters (DT). In the second case, he trains a surrogate decision tree with the exact same hyperparameters as the target model (DT+). Finally, in the third case, a multi-layer perceptron (MLP) is used as surrogate model. Both CF and DualCF were originally tested using the DiCE (Mothilal et al., 2020) counterfactual oracle, which provides heuristic-based (non-optimal) counterfactual explanations. To assess the impact of explanation optimality on the attack’s performance, and to ensure fair comparisons, we run CF and DualCF both using DiCE and using optimal counterfactual explanations computed with the OCEAN framework (Parmentier & Vidal, 2021).
@@
We hereafter highlight the key trends of these results, on two dimensions: the adversarial knowledge (regarding the target model’s architecture) and the type of counterfactual oracle.
@@
One first important trend that is consistent across both CF and DualCF, and for both decision trees and random forests, is that knowledge of the hypothesis class of the target model helps fitting a surrogate with high fidelity. Indeed, as can be observed in Figures 8, 9 and 10, the MLP surrogate always under-performs compared to the decision trees or random forests ones. Indeed, fitting a surrogate model of the same type is facilitated by the fact that the shapes of its decision boundary are the same as the target model, e.g., axis-parallel splits for tree-based models. Interestingly, knowledge of the hyperparameters of the target decision tree or random forests does not seem to help fitting the surrogate. Indeed, in most experiments, the surrogate sharing the same hypothesis class as the target model (i.e., DT or RF) and the surrogate sharing both the hypothesis class and the hyperparameters (i.e., DT+ or RF+) have very close performances. Furthermore, imposing the target model’s hyperparameters to the trained surrogate can even be counterproductive, as can be seen in Figures 8(a) and 8(b) for instance. In such cases, the fact that surrogate learning is more constrained due to the enforced hyperparameters (e.g., maximum depth) seems to slow its convergence towards very high fidelity values. This is consistent with previous findings: for instance, Aïvodji et al. (2020) observed that knowledge of the architecture of a target neural network did not provide a significant advantage to the model extraction attacker.
@@
The results in Figures 8, 9 and 10 suggest that the non-optimal counterfactual examples returned by DiCE helped fitting the extracted surrogate models better than the optimal ones provided by OCEAN. Indeed, for a fixed query budget and surrogate type, the performances of the extracted model are often better when fitted with DiCE counterfactuals than with OCEAN ones. Although some variations appear, this finding is generally verified for all the studied types of surrogates, for both decision trees and random forests target models, and for both the CF and DualCF extraction attacks. Intuitively, this can be attributed to a greater diversity in the non-optimal counterfactuals, which do not necessarily lie close to a decision boundary, unlike optimal ones. This also highlights a crucial insight: optimal counterfactuals only give an advantage to a model extraction attack if the attack is able to leverage the information it carries as a whole (including both the counterfactual example and its optimality) through a structured approach, as demonstrated by TRA.
@@ B.3 Test Set Fidelity Results
The results provided in Section 4.2 measure fidelity on a dataset uniformly sampled over the input space, which accurately quantifies how well the extracted models fit the decision boundaries of the target ones over the whole input space. Another approach consists in evaluating fidelity on a test set. In such cases, the results indicate how well the extracted models mimic the target one for examples drawn from the actual data distribution. We report in Table 2 the results of our extraction attacks against random forests. More precisely, for random forests of varying sizes, we report the average fidelity (measured on the uniformly sampled dataset or on a test set) achieved by all considered methods along with the required number of queries. For DualCF and CF, these values are arbitrarily fixed to allow their surrogates to converge towards (near) perfect fidelity. For our proposed TRA, functional equivalence is achieved using the reported number of queries, hence fidelity on both considered datasets is always $1.0$. For CF and DualCF, we report results for the random forest surrogate using default parameters, for both studied counterfactual oracles. Indeed, we observed in Section B.2 that DiCE counterfactuals led to better anytime performances (in terms of uniform dataset fidelity) than OCEAN ones within the CF model extraction attack, in most experiments. However, this is not always the case, with a few setups in which the difference between both approaches becomes very small or shifts in favor of the runs using OCEAN counterfactuals after sufficiently many iterations. This is the case on the COMPAS dataset (Figure 8(b)), and although the difference remains very subtle, it is also visible on the test set fidelity, illustrating the fact that the two values are often very well aligned.
Overall, the superiority of TRA is clear, both in terms of (uniform or test) fidelity and in terms of required numbers of queries, confirming the observation of Section 4.2.
@@
Dataset #Trees Nodes #Queries FU FTD #Queries FU FTD #Queries FU FTD #Queries FU FTD #Queries FU FTD
@@ B.4 Detailed Experimental Results
We hereafter report all the results of our main experiments over the six considered datasets.
First, Figure 11 provides the anytime performances (in terms of average surrogate fidelity as a function of the number of performed queries) of the four considered model extraction attacks when applied on decision tree target models. The findings highlighted in Section 4.2 are consistent across all considered datasets: TRA exhibits higher anytime fidelity than CF and DualCF for all query budgets, while also providing functional equivalence guarantees. While PathFinding also provides these guarantees, it necessitates orders of magnitudes more queries to fit its surrogate.
Figure 12 focuses on functionally equivalent model extraction attacks, and relates the number of queries they require to fully recover the target model to its size (quantified as its number of nodes). The logarithmic scale of the y-axis highlights that TRA usually requires orders of magnitudes less queries than PathFinding to entirely extract a given decision tree. Interestingly, this trend is more subtle when reconstructing large trees trained on the datasets with the highest numbers of features (i.e., Adult and SPerformance).
Finally, Figure 13 reports the number of counterfactual queries required by TRA to conduct a functionally equivalent extraction of random forests of various sizes, as a function of the total number of nodes to be recovered within the target forest. Importantly, as quantified through the performed power-law regression, the number of queries required by TRA to entirely extract the target forests grows sub-linearly – in $\Theta(\text{\#Nodes}^{0.38})$ – with the total number of nodes to be retrieved. This empirically demonstrates the very good scalability of TRA with respect to the size of the target random forests. Note that this behavior arises because large forests with many trees introduce redundancies, allowing the extracted model to be represented with perfect fidelity as a more compact decision tree (Vidal & Schiffer, 2020).

### v2

@@ Appendix A Proofs
    Each subspace contains subtrees $f_{n_{1}}$ and $f_{n_{2}}$ with $n_{1},n_{2}\leq n$ split levels, respectively. By the inductive hypothesis, the number of queries for each subtree is respectively $O\left(\prod_{\begin{subarray}{c}i=1\end{subarray}}^{m}(s^{(n_{1})}_{i}+1)\right)$ and $O\left(\prod_{\begin{subarray}{c}i=1\end{subarray}}^{m}(s^{(n_{2})}_{i}+1)\right)$.
@@
    $O\left(\prod_{\begin{subarray}{c}i=1\end{subarray}}^{m}(s^{(n+1)}_{i}+1)\right).$
Therefore, the query count for extracting $f_{n}$ is $O\left(\prod_{\begin{subarray}{c}i=1\end{subarray}}^{m}(s^{(n)}_{i}+1)\right)$. ∎
@@
$\max_{s^{(n)}_{1},\ldots,s^{(n)}_{m}}\prod_{i=1}^{m}(s^{(n)}_{i}+1)\quad\text{s.t.}\quad\sum_{i=1}^{m}s^{(n)}_{i}=n\quad\text{and}\quad s^{(n)}_{i}\geq 1\ \forall i\in\{1,\ldots,m\}.$
@@
$\max_{s^{(n)}_{1},\ldots,s^{(n)}_{m}}\sum_{i=1}^{m}\log(s^{(n)}_{i}+1).$
Applying the Karush-Kuhn-Tucker (KKT) conditions, we find that the maximum occurs when $s^{(n)}_{i}=\frac{n}{m}$ for all $i$. Substituting back, the worst-case complexity becomes $O\left(\left(\frac{n}{m}+1\right)^{m}\right)$. ∎
@@
$\forall 1\leq j\leq m,\quad\sum_{i=1}^{j-1}s_{i}+1\leq p\leq\sum_{i=1}^{j}s_{i},\quad\alpha_{p}<\alpha_{p+1}.$
@@
    In the best-case scenario, where each split level appears exactly once in the decision tree (i.e., there is no redundancy among the tree nodes), any algorithm, including the optimal one, would require at least $12n+$ 1 queries to reconstruct the tree:
    $Q_{opt}^{f}\geq 2n+1$ (2)
    This includes one query for discovering each split and additional queries to verify the leaf nodes.
    Conversely, in the worst-case scenario, such as a chessboard-like decision tree where splits are evenly distributed across multiple features, the TRA algorithm must explore all possible regions created by these splits. For a two-dimensional tree, this results in $s_{1}+s_{2}(s_{1}+1)+(s_{1}+1)(s_{2}+1)$ queries, where $s_{1}$ and $s_{2}$ are the number of splits along each feature. The $s_{1}$ splits along the first dimension are first detected, then the $s_{2}$ splits along the second dimension are re-discovered at every sub-division performed along the first dimension, and finally $(s_{1}+1)(s_{2}+1)$ queries are required to individually verify each sub-square (leaf node). Extending this to $m$ dimensions, the number of queries grows multiplicatively with the number of splits per feature, leading to:
@@
    $C_{TRA}^{(n,m)}=\text{sup}_{f\in\mathcal{F}}\left(\frac{Q_{TRA}^{f}}{Q_{opt}^{f}}\right)\leq\frac{\sum_{i=1}^{m}s_{i}\prod_{j=1}^{i-1}(s_{j}+1)+\prod_{j=1}^{m}(s_{j}+1)}{2n+1}$ (3)
@@
    Proof of Lower Bound: We hereafter build an adversarial example, i.e., one that maximizes the ratio of the number of queries that TRA must perform to extract the target decision tree, compared to what an optimal offline algorithm could achieve. This example therefore constitutes a (feasible) lower bound for the competitive ratio of TRA.
@@
    Consider the following adversarial example: for each $1\leq p\leq n$ and $1\leq j\leq m$, let the dimension that $\alpha_{p}$ splits on be $j$, and set
    $\alpha_{p}=\begin{cases}\frac{p}{(s_{1}+1)},&\text{if }j=1,\\ \frac{p-\sum_{i=1}^{j-1}s_{i}}{2(s_{j}+1)}+\frac{1}{2}+\epsilon,&\text{otherwise},\end{cases}$
    where $\epsilon>0$. This adversarial example ensures that within any hyper-rectangle defined by split level boundaries, for $j=1,...,m-1$ if there are splits in both the $j$-th and $j+1$-th dimensions, then there exists a split in the $j$-th dimension that is closer to the center of the hyper-rectangle than any split in the $j+1$-th dimension. As a consequence, TRA will always detect the splits of the $j$-th dimension before those of the $j+1$-th dimension. Therefore, the adversary can design a decision tree with a single branch (as illustrated in the right tree of Figure 6) that begins by splitting on the split levels in reverse (decreasing) order of dimensions (see a 2D example in Figure 6). For this specific example, TRA will require
@@
    $C_{TRA}^{(n,m)}\geq\frac{\sum_{i=1}^{m}s_{i}\prod_{j=1}^{i-1}(s_{j}+1)+\prod_{j=1}^{m}(s_{j}+1)}{2n+1}$ (4)
Hence, by (3) and (4), we have:
@@
Figure 6: Adversarial example for TRA (displayed for $m=2$ dimensions). The classes are $c_{1}$,$c_{2}$ and $c_{3}$. For simplicity, we denote $s_{1}=s^{(n)}_{1}$ and $s_{2}=s^{(n)}_{2}$. Here, the instance is dynamically built so that the number of queries required by TRA is $s_{2}(s_{1}+1)+s_{1}+(s_{1}+1)(s_{2}+1)$, whereas the optimal offline algorithm only needs $12n+$ 1, as shown in the right tree figure.
@@
We define a pure D&C-based algorithm as one that divides the input space (problem) into subspaces (sub-problems) based on counterfactual explanations and recursively continues this process within each subspace until no counterfactuals are found. This class of algorithms encompasses all types of querying strategies, such as selecting the geometrical center (as done by TRA), but also selecting the top-left corner, bottom-right corner, or a random point within the input space, among others.
@@
    $1+{Q}_{1}+{Q}_{2}=(2n+1)C_{TRA}^{(n,m)}$ (5)
    given that $s_{1}^{(1)}+s_{1}^{(2)}+1=s_{1}$. Hence, by (2) and (5), the best competitive ratio $C^{(n,m)}_{D\&C}$ achievable by any D&C-based algorithm satisfies:
@@ B.1 Experimental Setup Details
For surrogate models (used by the CF and DualCF attacks) that do not utilize the target model’s hyperparameters, we employ the default parameters provided by the scikit-learn Python library [Pedregosa et al., 2011]. Specifically for MLPs, we configure a scikit-learn MLP with two hidden layers, each consisting of 20 neurons, while keeping all other parameters at their default values.
@@
The anytime fidelity was calculated each 20 queries during all attacks execution, except for PathFinding which is not an anytime attack. Let $N$ denote the number of target models $f_{1},f_{2},\ldots,f_{N}$ to extract (i.e., over all considered experimental configurations and random seeds) and $\mathcal{D}$ a dataset with $n$ samples. The anytime fidelity of a given model extraction attack (at a given time step) over $\mathcal{D}$ is calculated as follows :
$\frac{1}{N}\sum_{i=1}^{N}\left(\frac{1}{n}\sum_{j=1}^{n}\mathds{1}_{\{f_{i}(x_{j})=\hat{f}_{i}(x_{j})\}}\right)$ (6)
where $\hat{f}_{1},\hat{f}_{2},\ldots,\hat{f}_{N}$ are the models extracted by the considered model extraction attack at the given time step.
@@ B.2 Configuration of Surrogate-Based Attacks
We report in Figure 8 (respectively, Figure 9) the anytime performance of the CF (respectively, DualCF) model extraction attack against decision tree models, for the three considered types of surrogates and the two counterfactual oracles, on all considered datasets. More precisely, as depicted in Section 4.1, we run these two attacks using three different assumptions on adversarial knowledge, namely the hypothesis class of the target model, its hyperparameters, and none of them. In the first case, the adversary trains a decision tree surrogate with default parameters (DT). In the second case, he trains a surrogate decision tree with the exact same hyperparameters as the target model (DT+). Finally, in the third case, a multi-layer perceptron (MLP) is used as surrogate model. Both CF and DualCF were originally tested using the DiCE [Mothilal et al., 2020] counterfactual oracle, which provides heuristic-based (non-optimal) counterfactual explanations. To assess the impact of explanation optimality on the attack’s performance, and to ensure fair comparisons, we run CF and DualCF both using DiCE and using optimal counterfactual explanations computed with the OCEAN framework [Parmentier and Vidal, 2021].
@@
We hereafter highlight the key trends of these results, focusing on the impact of two dimensions: the adversarial knowledge (regarding the target model’s architecture and hyperparameters) and the type of counterfactual oracle used.
@@
One first important trend that is consistent across both CF and DualCF, and for both decision trees and random forests, is that knowledge of the hypothesis class of the target model helps fitting a surrogate with high fidelity. Indeed, as can be observed in Figures 8, 9 and 10, the MLP surrogate always under-performs compared to the decision trees or random forests ones. Indeed, fitting a surrogate model of the same type is facilitated by the fact that the shapes of its decision boundaries are the same as the target model, e.g., axis-parallel splits for tree-based models. Interestingly, knowledge of the hyperparameters of the target decision tree or random forests does not seem to help fitting the surrogate. Indeed, in most experiments, the surrogate sharing the same hypothesis class as the target model (i.e., DT or RF) and the surrogate sharing both the hypothesis class and the hyperparameters (i.e., DT+ or RF+) have very close performances. Furthermore, imposing the target model’s hyperparameters to the trained surrogate can even be counterproductive, as can be seen in Figures 8(a) and 8(b) for instance. In such cases, the fact that surrogate learning is more constrained due to the enforced hyperparameters (e.g., maximum depth) seems to slow its convergence towards very high fidelity values. This is consistent with previous findings: for instance, Aïvodji et al. [2020] observed that knowledge of the architecture of a target neural network did not provide a significant advantage to the CF model extraction attack.
@@
The results in Figures 8, 9 and 10 suggest that the non-optimal counterfactual explanations returned by DiCE helped fitting the extracted surrogate models better than the optimal ones provided by OCEAN. Indeed, for a fixed query budget and surrogate type, the performances of the extracted model are often better when fitted with DiCE counterfactuals than with OCEAN ones. Although some variations appear, this finding is generally verified for all the studied types of surrogates, for both decision trees and random forests target models, and for both the CF and DualCF extraction attacks. Intuitively, this can be attributed to a greater diversity in the non-optimal counterfactuals, which do not necessarily lie close to a decision boundary, unlike optimal ones. This also highlights a crucial insight: optimal counterfactuals only give an advantage to a model extraction attack if the attack is able to leverage the information it carries as a whole (including both the counterfactual example and its optimality) through a structured approach, as demonstrated by TRA.
@@ B.3 Test Set Fidelity Results
As mentioned in Section 4.1, the results provided in Section 4.2 measure fidelity on a dataset uniformly sampled over the input space, which accurately quantifies how well the extracted models fit the decision boundaries of the target ones over the whole input space. Another approach consists in evaluating fidelity on a test set. In such cases, the results indicate how well the extracted models mimic the target ones for examples drawn from the actual data distribution. We report in Table 2 the results of our extraction attacks against random forests. More precisely, for random forests of varying sizes, we report the average fidelity (measured on the uniformly sampled dataset or on a test set) achieved by all considered methods along with the required number of queries. For DualCF and CF, these values are arbitrarily fixed to allow their surrogates to converge towards (near) perfect fidelity. For our proposed TRA, functional equivalence is achieved using the reported number of queries, hence fidelity on both considered datasets is always $1.0$. For CF and DualCF, we report results for the random forest surrogate using default parameters, for both studied counterfactual oracles. Indeed, we observed in Section B.2 that DiCE counterfactuals led to better anytime performances (in terms of uniform dataset fidelity) than OCEAN ones within the CF model extraction attack, in most experiments. However, this is not always the case, with a few setups in which the difference between both approaches becomes very small or shifts in favor of the runs using OCEAN counterfactuals after sufficiently many iterations. This is the case on the COMPAS dataset (Figure 8(b)), and although the difference remains very subtle, it is also visible on the test set fidelity, illustrating the fact that the two values are often very well aligned.
Overall, the superiority of TRA is clear, both in terms of (uniform or test) fidelity and in terms of required numbers of queries, confirming the observations of Section 4.2.
@@
Dataset #Trees #Nodes #Queries FU FTD #Queries FU FTD #Queries FU FTD #Queries FU FTD #Queries FU FTD
@@ B.4 Detailed Experimental Results
We hereafter report all the results of our main experiments over the five considered datasets.
First, Figure 11 provides the anytime performances (in terms of average surrogate fidelity as a function of the number of performed queries) of the four considered model extraction attacks when applied on decision tree target models. The findings highlighted in Section 4.2 (in particular, Result 1) are consistent across all considered datasets: TRA exhibits higher anytime fidelity than CF and DualCF for all query budgets, while also providing functional equivalence guarantees. While PathFinding also provides these guarantees, it necessitates orders of magnitudes more queries to fit its surrogate.
Figure 12 focuses on functionally equivalent model extraction attacks, and relates the number of queries they require to fully recover the target model to its size (quantified as its number of nodes). The logarithmic scale of the y-axis highlights that TRA usually requires orders of magnitudes less queries than PathFinding to entirely extract a given decision tree, confirming our Result 2 (Section 4.2). Interestingly, this trend is more subtle when reconstructing large trees trained on the datasets with the highest numbers of features (i.e., Adult and SPerformance).
Finally, Figure 13 reports the number of counterfactual queries required by TRA to conduct a functionally equivalent extraction of random forests of various sizes, as a function of the total number of nodes to be recovered within the target forest. Importantly, as quantified through the performed power-law regression, the number of queries required by TRA to entirely extract the target forests grows sub-linearly – in $\Theta(\text{\#Nodes}^{0.38})$ – with the total number of nodes to be retrieved. This empirically demonstrates the very good scalability of TRA with respect to the size of the target random forests, consistent with Result 3 (Section 4.2). Note that this behavior arises because large forests with many trees introduce redundancies, allowing the extracted model to be represented with perfect fidelity as a more compact decision tree [Vidal and Schiffer, 2020].
@@
Appendix C Performances of TRA when used with Locally Optimal Counterfactuals
In the main body of the paper (Section 4.2) we employed the OCEAN counterfactual oracle for our experiments using TRA. It supplies globally optimal counterfactual explanations. While such global optimality is sufficient for the TRA attack to succeed, it is not necessary. Indeed, as pointed out in Section 3.2, TRA requires counterfactuals that are locally optimal – i.e., positioned on (or infinitesimally close to) the decision boundary of the explained model $f$. To illustrate this distinction, we replicate all experiments, following exactly the same setup described in Section 4.1, using locally optimal counterfactuals.
The DiCE counterfactual oracle, used in our experiments for the CF and DualCF model extraction attacks, could be a good candidate. However, it frequently fails to return valid counterfactuals when restricted to a given subspace of the input space – making it difficult to use within TRA. This suggests that the counterfactuals found by DiCE often lie in different decision regions of the target model, rather far from the actual query. When the counterfactuals are used to provide recourse information, this can be problematic – for instance, by suggesting that a credit applicant must exert significantly more effort than actually necessary to improve their application. For these reasons, in our experiments using TRA, we rather consider a lightweight heuristic (Algorithm 2) producing locally optimal counterfactuals.
Algorithm 2 A simple heuristic to find locally optimal counterfactual explanations
1:  Input: Query $x$, model $f$, input space $\mathcal{E}$, training data $\mathcal{D}_{T}$, and maximum number of iterations for the uniform sampling process $S$.
2:  Return: A locally optimal counterfactual explanation $x^{\prime}\in\mathcal{E}$ if it exists.
3:  $\mathcal{D}_{\mathcal{E}}\leftarrow\mathcal{D}_{T}\cap\mathcal{E}$gets all the data points that are in the desired input space $\mathcal{E}$}
4:  for $x^{\prime}\in\mathcal{D}_{\mathcal{E}}$  do
5:     if  $f(x^{\prime})\neq f(x)$  then
6:        return  $linesearch(x,x^{\prime})$$x^{\prime}$ is a counterfactual, we refine it via line search towards $x$}
7:     end if
8:  end for
9:  for  $0\leq i\leq S$ do
10:     $x^{\prime}\leftarrow sample(\mathcal{E})$sample a point $x^{\prime}$ uniformly within $\mathcal{E}$}
11:     if  $f(x^{\prime})\neq f(x)$  then
12:        return  $linesearch(x,x^{\prime})$$x^{\prime}$ is a counterfactual, we refine it via line search towards $x$}
13:     end if
14:  end for
15:  return  $None$no counterfactual found}
More precisely, Algorithm 2 searches in the training dataset and then refines the first valid counterfactual it finds via a one-dimensional line search (line 6). If no counterfactual is found, it tries another strategy by sampling a predefined number of points uniformly in $\mathcal{E}$. In our experiments, the maximum number of such sampled points is fixed to $S=1000$. If a valid counterfactual is found, it is refined via line search and returned (line 12). Because the search is stochastic and local, it may fail to return a counterfactual even when one exists, yet it satisfies local optimality whenever a solution is found.
Figures 14 and 15 contrast the performance of TRA to extract decision trees, using OCEAN (globally optimal counterfactuals) versus Algorithm 2 (simpler algorithm, producing locally optimal counterfactuals) as counterfactual oracle. The observed trends confirm show that TRA remains effective as long as the returned counterfactuals are locally, though not necessarily globally, optimal – thereby confirming Result 4 (Section 4.2). Interestingly, we observe in Figure 14 that in the early iterations, locally optimal counterfactuals may lead to better anytime fidelity values – which is likely the case due to their heuristic nature, enhancing diversity. We also observe a pathological case in Figure 14(a), with the Adult dataset. In this particular experiment, when used with Algorithm 2 as counterfactual oracle, TRA converges close to near-perfect – but not exact – fidelity. This is due to the fact that Algorithm 2 fails to find feasible counterfactuals within the desired region, even if one exists. Indeed, even if the counterfactuals provided by this simple oracle satisfy local optimality whenever one is found, there is no guarantee if none is found that none actually exists. Performing a finer uniform sampling (via raising the number of sampled points $S$) could fix this issue. Nevertheless, the performances of TRA with this weak oracle remain very robust, as evidenced through our large set of experiments.
The performance of TRA to extract random forests target models, with either the OCEAN or Algorithm 2 counterfactual oracles, are provided in Table 3 and display the same trends. TRA is able to achieve perfect fidelity for both counterfactual oracles, and there is no significant difference in the required number of queries (which remains very low compared to the size of the reconstructed forests, as discussed in Section 4.2).
(a) Adult dataset
(b) COMPAS dataset
(c) Credit Card dataset
(d) German Credit dataset
(e) Student Performance dataset
Figure 14: Anytime performance of TRA with either OCEAN or a simpler heuristic counterfactual oracle (Algorithm 2) to extract decision trees. We report results for all datasets.
(a) Adult dataset
(b) COMPAS dataset
(c) Credit Card dataset
(d) German Credit dataset
(e) Student Performance dataset
Figure 15: Performance of TRA with either OCEAN or a simpler heuristic counterfactual oracle (Algorithm 2) to achieve functionally equivalent model extraction attacks against decision trees. We report results for all datasets where each point represents the number of queries needed to fully reconstruct the trees.
Table 3: Summary of our model extraction experiments against random forests, on the COMPAS dataset using both OCEAN and Heuristic. FU and FTD denote respectively the Fidelity over the Uniform and Test Data.
TRA
OCEAN Heuristic
Dataset #Trees Nodes #Queries FU FTD #Queries FU FTD
COMPAS 5 486.60 73.60 1.00 1.00 74.20 1.00 1.00
25 4569.00 138.80 1.00 1.00 140.00 1.00 1.00
50 9151.20 147.60 1.00 1.00 149.20 1.00 1.00
75 7317.00 95.20 1.00 1.00 95.20 1.00 1.00
100 18369.20 129.60 1.00 1.00 130.40 1.00 1.00
Appendix D Broader Impact Statement
As a result, MLaaS platforms increasingly integrate such explainability tools into their APIs. While these explanations enhance user trust, they also expose a new attack surface to malicious entities by revealing additional model information. In this work, we theoretically and empirically demonstrate that locally optimal counterfactual explanations of decision trees and tree ensembles can be exploited to conduct efficient model extraction attacks. These results highlight the critical tension between transparency and the protection of model integrity and intellectual property. By identifying and quantifying these vulnerabilities, we highlight the risks of releasing model explanations without a thorough security assessment. Our research establishes a benchmark for evaluating method safety, advocating for the development of privacy-preserving approaches to explainability.